OpenStudy (anonymous):

Help please! Will give medal!(: Simplify : (y + 3) * 2 = (-1) * ( x - 4)

OpenStudy (kropot72):

To simplify the left hand side, multiply each term inside the brackets by 2. Can you do that?

OpenStudy (anonymous):

I think so but, i'm not sure about the right side, do I multiply the (-1) or (x - 4) by 2??

OpenStudy (kropot72):

You don't need to apply 2 to the right hand side. Can we treat the right hand side separately?

OpenStudy (anonymous):

yes, s o would the left side be , 2y + 6?

OpenStudy (kropot72):

Good work! Now for the right hand side: ignore the brackets around -1, and multiply each term inside the right hand brackets by -1.

OpenStudy (anonymous):

so it would be -1 and -4

OpenStudy (anonymous):

2y+6=4-x

OpenStudy (kropot72):

@cody_123 Do you want to carry on?

OpenStudy (kropot72):

(-1) * ( x - 4) = -1(x - 4) = -x +4 = 4 - x

OpenStudy (kropot72):

So now we have simplified to: 2y + 6 = 4 - x The next step is to add x to the left hand and the right hand sides. Can you do that?

OpenStudy (anonymous):

2y + x ?

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

2y + 6 + x = 4 -x + x This can be simplified and rearranged to: x + 2y + 6 = 4 Now we need to subtract 6 from both left and right hand sides: x + 2y + 6 - 6 = 4 - 6 ............(1) Can you now simplify (1) to get the answer?

OpenStudy (anonymous):

2y + x = 5 ?

OpenStudy (kropot72):

x + 2y + 6 - 6 = 4 - 6 ............(1) The left hand side becomes x + 2y, the reason being that 6 - 6 = 0. So we now have: x + 2y = 4 - 6 ...............(2) Can you simplify the right hand side?